### v1

# A Graph Theoretic Approach for Exploring the Relationship between EV Adoption and Charging Infrastructure Growth

Abstract
The increasing global demand for conventional energy has led to significant challenges, particularly due to rising CO2 emissions and the depletion of natural resources. In the U.S., light-duty vehicles contribute significantly to transportation sector emissions, prompting a global shift toward electrified vehicles (EVs). Among the challenges that thwart the widespread adoption of EVs is the insufficient charging infrastructure (CI). This study focuses on exploring the complex relationship between EV adoption and CI growth. Employing a graph theoretic approach, we propose a graph model to analyze correlations between EV adoption and CI growth across 137 counties in six states. We examine how different time granularities impact these correlations in two distinct scenarios: Early Adoption and Late Adoption. Further, we conduct causality tests to assess the directional relationship between EV adoption and CI growth in both scenarios. Our main findings reveal that analysis using lower levels of time granularity result in more homogeneous clusters, with notable differences between clusters in EV adoption and those in CI growth. Additionally, we identify causal relationships between EV adoption and CI growth in 137 counties, and show that causality is observed more frequently in Early Adoption scenarios than in Late Adoption ones. However, the causal effects in Early Adoption are slower than those in Late Adoption.
1 Introduction
As global populations grow, the demand for conventional energy continues to rise, leading to increased CO2 emissions and depletion of finite natural resources [1]. In the U.S., transportation contributes nearly 29% of carbon emissions, with light-duty vehicles responsible for 59% of these. This highlights the urgency of transitioning to sustainable energy sources [2, 3].
Achieving zero carbon emissions in light-duty transportation is crucial, and electric vehicles (EVs) offer a promising solution, as they produce zero tailpipe emissions [4]. However, widespread EV adoption faces obstacles, such as insufficient charging infrastructure, high upfront costs, and concerns over limited range [5]. Government interventions, including subsidies for charging infrastructure (CI) and tax credits for EV purchases, aim to address these barriers. Despite these efforts, the relationship between EV adoption and CI development remains complex: does the availability of charging stations drive EV adoption, or does widespread EV use lead to more charging infrastructure? Furthermore, the impacts of one on the other may emerge over different time granularities.
This study explores the relationship between EV adoption (represented by the EV growth rate) and CI growth using a graph-theoretic approach to propose a graph model for analyzing correlations in EV adoption across 137 counties in six U.S. states. In the proposed graph model, nodes represent counties, and edges represent correlations between them. Our analysis constructs a graph model to capture correlations at various time granularities: monthly, bi-monthly, quarterly, and bi-annual. After constructing a graph for each time granularity, we apply the Louvain method, a community detection technique, to cluster counties into different groups. Similarly, we perform the same analysis on correlations in CI growth across these counties. We then examine the relationship between EV adoption and CI growth considering different time lags to explore Early Adoption scenarios (where EV adoption leads CI growth) and Late Adoption scenarios (where CI growth leads EV adoption). Finally, we analyze causal relationships between EV adoption and CI growth by considering various combinations of time granularities and lags to better understand these complex interactions.
Firstly, our analysis reveals that, within EV adoption networks, counties from the same state tend to cluster together, exhibiting greater homogeneity at finer time granularities. However, this pattern is not observed in CI growth networks. Secondly, the correlations between EV and CI networks in both adoption scenarios, Early Adoption and Late Adoption, are consistently weakly positive, with Late Adoption correlations being slightly higher than those in Early Adoption. Thirdly, 115 out of 137 counties exhibit causal relationships between EV adoption and CI growth, analyzed using various combinations of time granularities, time lags, and the two adoption scenarios. Notably, causality is observed more frequently in Early Adoption scenarios than in Late Adoption ones. However, the causal effects in Early Adoption are slower than those in Late Adoption, which can be attributed to the significantly higher costs of establishing new CIs compared to adopting new EVs.
The remainder of this paper is organized as follows. In Section 2, we review the related work to the EV adoptions and CI growth. Section 3 discusses our approach for employing a graph-theoretic approach to build the EV adoption’s correlation networks, and the CI growth’s correlation networks, and for the causality tests. Finally, Section 4 discusses the results followed by the conclusion and future work in Section 5.
2 Literature Review
In this section, we review existing research on the complex relationship between EV adoption and CI growth. In addition, we discuss studies that have utilized a graph model in the context of CI optimization and identify gaps in the literature that our study aims to address. This review sets the stage for understanding the unique contributions of our research to the field.
2.1 Impact of Charging Infrastructure and Other Factors on EV Adoption
Ledna et al. investigated the effect of various support policies for infrastructure investment and vehicle purchases on EV adoption in California [6]. They analyzed three policy scenarios: no support, support for either infrastructure or vehicle purchases, and combined support for both. These scenarios were assessed under conservative and optimistic technological advancement assumptions. Their findings indicated that the combined policy best supported vehicle sales and reduced CO2 emissions, while support focused on infrastructure alone proved more effective in conservative scenarios.
In another study, Chen et al. proposed a game-theoretic model to understand the interactions between the government, technology firms, and consumers in the EV market [5]. Their objective was to identify optimal government incentives to maximize market penetration while enhancing consumer surplus and social welfare. Their research revealed that, under a high-budget incentive framework, usage subsidies best maximized welfare, while investment subsidies were optimal for market penetration.
White et al. examined the relationship between the intent to adopt EVs and charging infrastructure through mediation analyses in three U.S. metropolitan areas [7]. They evaluated the indirect effects of Range Anxiety, Perceived Mobility Restriction (PMR), and Subjective Norms on adoption intent. Their results demonstrated that Subjective Norms were the strongest mediator, while Range Anxiety was marginally significant, and PMR showed no significant effect.
Khan et al. hypothesized that socio-demographic features and highway routes influence the presence of charging stations at the zip-code level [8]. They compared zip codes with and without stations using a t-test, finding that the presence of highways, income, and education levels positively correlated with charging station distribution.
At a country level, in Australia, Broadbent et al. proposed a framework to estimate the number of EVs in parking lots in Melbourne and transform them into EV-enabled spaces [9]. In Germany, Illmann et al. analyzed monthly data from 2011 to 2017 and found a long-term relationship between registered EVs and charging stations [10]. In France, Haidar et al. discovered that different incentives impacted hybrid and fully electric vehicles differently [11].
2.2 Application of Graph Model to Charging Infrastructure
The graph model is a powerful tool for analyzing relationships between various entities. A correlation network, where nodes represent entities or variables, and edges indicate correlation strength, is widely used in diverse fields, including bioinformatics, engineering, and social sciences.
The application of the graph model to the charging infrastructure has focused primarily on optimizing the distribution of charging stations [12]. For example, Gagarin et al. applied the dominating set concept to create a network of charging stations reachable within a specific threshold [13]. Similarly, Arkin et al. utilized Unit Disk Graphs to determine the minimum independent set required to cover an area with charging stations [14]. Moreover, studies in [15, 16, 17] applied the concept of the hitting set to identify the optimal number and locations for charging stations. Finally, Altundogan et al. combined genetic algorithms and shortest path optimization to determine the best locations for a pre-determined number of charging stations [18].
2.3 Gaps Found in the Literature
As we discussed the literature in sections 2.1 and 2.2, we noticed two gaps in the literature which the current study addresses:
  1. 1.
    Most of the existing literature examines EV adoption and its relationship with other factors in one to three cities, resulting in limited data diversity. This study, however, analyzes 137 counties across six U.S. states, enhancing the analysis and demonstrating the role of states in clustering counties within the correlation networks.
  2. 2.
    The use of graph model has mostly been limited to optimizing the placement of charging stations within specific areas. In contrast, this study contributes by applying graph model to explore patterns in both EV adoption and CI growth networks under two scenarios of Early Adoption and Late Adoption , build a correlation network, and cluster counties accordingly.
  3. 3.
    Our analysis examines two scenarios: Early Adoption and Late Adoption. We use different time granularities and lags to capture complex correlations between counties in both EV adoption and CI growth, as well as to identify potential causal relationships between EV adoption and CI growth within each county.
3 Methodology
In this section, we outline the approach used to analyze the relationship between EV adoption and charging infrastructure (CI) growth. It starts by describing the data in Section 3.1, using graph model to build correlation networks, introducing the two cases of adoption in Section 3.3, setting different ime granularities in Section 3.4, and setting two causality tests in Section 3.5
3.1 Data Collection
3.1.1 Electric Vehicles (EV) Data
Table 1: Number of Completed Counties by State
No State # of Completed Counties
1 Colorado 20
2 Minnesota 3
3 Montana 2
4 New York 48
6 Texas 30
7 Virginia 34
8 Total 137
The data used for electric vehicle (EV) registrations was obtained from Atlas Hub[19]; which works directly with states to provide comprehensive temporal data across various U.S. states, primarily at the zip code and county levels. For this study, we selected states offering consistent data from 2018 to 2023 at the monthly basis; so counties that had 12 months of data for each year in the selected period were included in the final dataset. Consequently, this dataset consists of EV data from 137 counties across six states; we aggregate the data at various time granularities for analyzing the data at different time scales. Table 1 provides a summary of the states and the number of counties from each.
3.1.2 Charging Infrastructure (CI) Data
Data on public and private charging stations were collected from the Alternative Fueling Station Locator, which, at the time of this research, included a total of 67,540 stations [20]. To ensure consistency, the dataset was filtered to include only the counties represented in the EV data. More importantly, the CI for a county is represented by the number of chargers rather than the number of charging stations, as it more accurately reflects the infrastructure’s capacity. Consequently, our analysis investigates the relationships between EV adoption and different CI levels, including Level 2 chargers, Direct Current (DC) chargers, and a combined level that includes Level 1, Level 2, and DC chargers, referred to as “ALL.” Additionally, each CI level for a county is aggregated into various time granularities, consistent with the approach used for the EV adoption data.
3.2 Proposed Graph Theoretic approach: correlation network
Our methodology applies modularity optimization to determine the optimal correlation threshold for each correlation network. The threshold that yields the highest modularity is selected as the most suitable for that network. This approach ensures that the network structure, defined by the retained correlations, reveals the most meaningful and interpretable relationships. Subsequently, the Louvain method is employed to detect communities (clusters) within the network.
3.3 Early versus Late Adoption
We define two adoptions scenarios of Early Adoption and Late Adoption to examine the relationship between EV adoption and CI growth using different time granularities.
3.3.1 Early Adoption
In this scenario, we assume that EVs are adopted before sufficient CI is available. In other words, EV adopters take the risk of using EVs despite a lack of supporting infrastructure, expecting their adoption to prompt further CI development. To analyze this, we apply a one-year lag, where EV adoption data spans 2018 to 2022 while CI growth data covers 2019 to 2023. This means a given year in the EV data (e.g., 2020) is aligned with the following year in the CI data (e.g., 2021). In this scenario, we investigate whether correlations exist between EV adoption and each level of CI, accounting for the one-year lag between the two datasets.
3.3.2 Late Adoption
This scenario assumes that CI is established before widespread EV adoption, with the expectation that sufficient CI will encourage more consumers to adopt EVs. Similar to the Early Adoption scenario, we apply a one-year lag: CI growth data spans 2018 to 2022, while EV adoption data covers 2019 to 2023. Here, CI data for a given year is aligned with EV data from the following year (e.g., 2018 CI data is aligned with 2019 EV data.)
3.4 Different Time Granularities
The relationship between EV adoption and CI growth is complex; therefore, our analysis considers using various time granularities to capture hidden relationships that may emerge at different time scales. To this end, we divide the dataset into multiple time granularities: monthly, bi-monthly, quarterly, and bi-annual periods. For each granularity, we calculate the growth rate from one period to the next. For instance, in the monthly granularity, we compute the growth rate from January to February, February to March, and so on. For the quarterly granularity, we calculate growth from the first quarter to the second quarter, and so forth.
Through this analysis, we examine the correlation networks of EV adoption and the growth of various CI levels, observing these correlations across different time scales to uncover potential relationships between EV adoption and CI growth. Specifically, we investigate whether clusters observed at finer granularities persist at coarser levels or whether new correlations between counties emerge, thereby shedding light on the temporal dynamics of EV adoption and CI expansion.
3.5 Causality Relationships between EVs and CIs
To explore the causal relationship between EV adoption and CI growth, we formulate two hypotheses:
  • Early Adoption Scenario: In this case, we test whether EV adoption leads to the development of CI, expressed as Null Hypothesis 1: "EV adoption does not cause CI growth." We evaluate this hypothesis using the Granger causality test.
  • Late Adoption Scenario: Here, we test whether CI growth drives EV adoption, expressed as Null Hypothesis 2: "CI growth does not cause EV adoption." We evaluate this hypothesis using the Granger causality test.
The Granger causality test is applied with an alpha level of 0.05, testing whether past values of one variable can predict future values of another. Additionally, we use multiple time lags—half-year, one-year, two-year, and three-year—to capture potential delayed effects of one on the other. These tests are conducted for different time granularity (monthly, bi-monthly, quarterly, and bi-annual), allowing us to identify causality across different time frames and adoption scenarios.
4 Results
In this section, we present the outcomes of our analysis, including the correlations for both EV adoption and CI growth using Graph Theory. We examine the two cases: Early Adoption (Section 4.1) and Late Adoption (Section 4.2) using different time granularity. Additionally, we present the results of our analysis of causality relationships between EV adoption and CI growth in these two cases in (Section 4.3).
A: Correlation Network for Monthly Data, Threshold: 0.7005
B: Correlation Network for Bi-monthly Data, Threshold: 0.8015
C: Correlation Network for Quarterly Data, Threshold: 0.8379
D: Correlation Network for Bi-annually Data, Threshold: 0.8394
Figure 1: Correlation Networks of EV Adoption (2018-2022) used in Early Adoption Scenario Using Different Time Granularities
A: Correlation Network for Monthly Data, Threshold: 0.7000
B: Correlation Network for Bi-monthly Data, Threshold: 0.7848
C: Correlation Network for Quarterly Data, Threshold: 0.8439
D: Correlation Network for Bi-annually Data, Threshold: 0.8500
Figure 2: Correlation Networks of ALL type of Chargers growth (2019-2023) used in Early Adoption Scenario Using Different Time Granularities
4.1 Early Adoption of EV
Graph 1 illustrates the correlation networks between counties for their EV adoptions (2018-2022) using different time granularity in the scenario of Early Adoption. The first notable finding is that lower levels of granularity produce more homogeneous clusters. For example, on a monthly basis, we observe that each cluster comprises counties from the same state. However, as the seasonal span increases, this homogeneity decreases, though it remains noticeable. Moreover, counties within the same cluster and state tend to be geographically close to each other. This raises the question of whether geographical proximity influences EV adoption—whether areas that are close to one another are more likely to adopt EVs if one area has already adopted them extensively (Contagion of EV Adoption.)
Similarly, Graph 2 illustrates the correlation between counties regarding the “ALL” level of CI (2019-2023) in the scenario of Early Adoption Scenario. Due to space limitations, and since the findings for “ALL” level chargers apply to other levels of CI, we do not show the other levels. In general, the clusters identified in CI do not exhibit the same patterns seen in EV adoption. The CI clusters are far more heterogeneous in terms of the states their counties belong to. The discrepancies between the EV and CI correlation patterns highlight the complexity of the relationship between EV adoption and CI development.
Lastly, Table 2 shows the correlation between EV adoption and various CI levels (Level 2 chargers, DC chargers, and ALL) in the scenario of Early adoption . We observe that the correlations are consistently weakly positive. Moreover, lower levels of granularity tend to have slightly stronger correlations than higher levels.
Table 2: Correlation between the EV adoption correlation and the correlation of various charging infrastructure types (DC chargers, Level 2 chargers, and ALL chargers) for Early Adoption Scenario, across different levels of temporal granularity.
Granularity EV vs DC EV vs Level 2 EV vs ALL
Monthly 0.1772 0.2205 0.2305
Bi-monthly 0.0972 0.1345 0.1416
Quarterly 0.0715 0.0624 0.0543
Bi-annually 0.0123 0.0366 0.0214
A: Correlation Network for Monthly Data, Threshold: 0.7020
B: Correlation Network for Bi-monthly Data, Threshold: 0.8015
C: Correlation Network for Quarterly Data, Threshold: 0.8379
D: Correlation Network for Bi-annually Data, Threshold: 0.8394
Figure 3: Correlation Networks of EV Adoption (2019-2023) used in Late Adoption Scenario Using Different Time Granularities
A: Correlation Network for Monthly Data, Threshold: 0.7110
B: Correlation Network for Bi-monthly Data, Threshold: 0.7848
C: Correlation Network for Quarterly Data, Threshold: 0.8439
D: Correlation Network for Bi-annually Data, Threshold: 0.8500
Figure 4: Correlation Networks of ALL type of Chargers growth (2018-2022) used in Late Adoption Scenario Using Different Time Granularities
4.2 Late Adoption of EV
Graph 3 illustrates the scenario of Late Adoption for the EV adoption (2019-2023). A similar pattern to Early Adoption scenario is observed, where lower levels of granularity show more homogeneity. However, there are noticeable differences between the clusters in the scenarios of Early and Late Adoption. One key difference is that the Early Adoption networks tend to be slightly denser than the Late Adoption networks. This could be due to the sparser nature of the early adoption data (2018-2022), which results in a higher number of correlations between counties.
For the correlations between counties regrading CI growth in the Late Adoption scenario, we only show the correlation for the “ALL” level of CI for simplicity in Graph 4. Analogously, the findings regarding CI promotion in the Late Adoption scenario are similar to those in Early Adoption, where the EV and CI correlation patterns differ significantly.
Lastly, Table 3 shows the correlation between EV adoption and various CI levels (Level 2 chargers, DC chargers, and ALL) in the Late Adoption scenario . We observe that the correlations are consistently weakly positive. Moreover, lower levels of time granularity tend to show slightly stronger correlations. Additionally, the correlation between EV adoption and CI correlation in the Late Adoption scenario is consistently higher than in the Early Adoption scenario across all levels of time granularity. This supports the conclusion that providing adequate charging infrastructure is crucial to promoting the use of EVs.
Table 3: Correlation between the EV adoption correlation and the correlation of various charging infrastructure types (DC chargers, Level 2 chargers, and ALL chargers) for late adoption, across different levels of temporal granularity.
Granularity EV vs DC EV vs Level 2 EV vs ALL
Monthly 0.1415 0.2529 0.2667
Bi-monthly 0.1806 0.1347 0.1824
Quarterly 0.0897 0.0737 0.0619
Bi-annually 0.0263 0.0368 0.0203
4.3 Causal Relationships Between EV Adoption and Charging Infrastructure
Table 4: The number of Counties Rejecting the Null hypothesis of the Causal Relationships in the Early Adoption Where the "EV" Causes the "ALL".
[…]
monthly 12 20 16 13 49
Bi-monthly 12 16 18 14 52
quarterly 17 19 17 15 57
Bi-annual 9 12 6 15 30
Unique 28 38 45 33 93
Table 5: The number of Counties Rejecting the Null hypothesis of the Causal Relationships in the Late Adoption Where the "ALL" Causes the "EV".
[…]
monthly 18 11 9 9 37
Bi-monthly 23 9 4 12 39
quarterly 15 6 4 8 27
Bi-annual 17 8 7 8 35
Unique 44 26 20 26 77
Table 6: The number of Counties Rejecting the Null hypothesis of the Causal Relationships in the Late Adoption Where the "ALL" Causes the "EV".
State EV->ALL ALL->EV Unique
New York 36 26 40
Texas 15 23 26
Virginia 23 16 27
Colorado 16 8 18
Minnesota 2 2 2
Montana 1 2 2
Unique 93 77 115
We conducted a Granger causality test to uncover the causal relationships between EV adoption and CI growth. In this analysis, we explore whether EV adoption causes CI growth or vice versa, as in the Early Adoption and Late Adoption scenarios, respectively. For simplicity, we focus on one level of CI, namely "ALL," which represents all types of chargers combined. Understanding the effects of specific levels, particularly DC chargers, on EV adoption is important and will be addressed in future work. In our causality analysis, we use different time granularities, including monthly, bi-monthly, quarterly, and bi-annual periods. Moreover, we test various lags between EV adoption and CI growth, including half-year, one-year, two-year, and three-year lags, to uncover potential relationships. Hence, each combination of a time granularity, a time lag, and an adoption scenario is tested for each county in our analysis. The number of counties rejecting the null hypotheses for Early Adoption and Late Adoption as explained in Section 4.3, is detailed in this section.
Table 4 presents the causal relationships captured in the Early Adoption scenario, where EV adoption causes CI "ALL" growth. Each entry in Table 4 represents a combination of a time granularity and time lag, indicating the number of counties with significant evidence of causal relationships. For instance, when using quarterly granularity and a two-year lag, 17 counties show such causal relationships. To account for counties appearing repeatedly across different combinations, the "Unique" column represents the unique counties showing causal relationships for a specific time granularity across different time lags, while the "Unique" row indicates the unique counties for a time lag across different granularities. Similarly, Table 5 presents the causal relationships captured in the Late Adoption scenario, where CI "ALL" growth causes EV adoption.
From Tables 4 and 5, we observe that the number of unique counties in the Early Adoption scenario (93 out of 137) is larger than in the Late Adoption scenario (77 out of 137). This indicates that EV adoption tends to cause CI growth more frequently than CI growth causes EV adoption. Consequently, EV adoption in an area before CI development is more critical than investing in CI before EV adoption. Thus, focusing on incentivizing EV purchases is a significant step for widespread adoption. Nevertheless, the demonstration of CI growth causing EV adoption in several counties highlights the complex and bidirectional nature of the relationship between EVs and CI.
Secondly, the causal effect of EV adoption on CI growth is clearly slower than the reverse. Specifically, the "Unique" row in Table 4 shows that the peak number of counties (45 counties) exhibiting CI growth caused by EV adoption occurs at a two-year lag. In contrast, the causal effect of CI growth on EV adoption peaks at a half-year lag (44 counties) as shown in Table 5. This finding is intuitive, as the slower causal effect of EV adoption on CI growth is likely due to the high costs associated with investing in CI infrastructure, such as charging stations. On the other hand, the causal effect of CI growth on EV adoption is faster, as purchasing an EV is an individual decision and typically less costly than establishing CI.
Overall, Table 6 shows that 115 unique counties out of 137 (approximately 84% of the data) exhibit significant causal relationships between EV adoption and CI growth using various time granularities, time lags, and adoption scenarios. Furthermore, Table 6 reveals that three states have a higher number of counties where EV adoption causes CI "ALL" growth, two states have more counties where CI "ALL" growth causes EV adoption, and one state exhibits an equal number of counties for both causal relationships.
4.4 Outlier Counties in EV Adoption Clusters
Focusing on the reasons behind the appearance of counties from different states within a cluster dominated by another state requires a more in-depth extension of this research. For instance, in the early adoption scenario using bi-monthly data, Danville City in Virginia appears in a cluster primarily composed of Texas counties. In this specific case, we identified shared factors such as rural influence, moderate commute distances, limited access to public transit, and lower median income, which may help explain similar patterns of EV adoption.
Conducting deeper analyses of these uncommon county appearances, both at the cluster level and across different adoption scenarios (Early vs. Late Adoption) can offer more nuanced insights into the factors driving EV adoption. Exploring these outliers may reveal underlying conditions that are not necessarily tied to geography or state policy but rather to shared structural or socioeconomic characteristics.
5 Conclusion and Future Directions
In this study, we examined the relationship between EV adoption and charging infrastructure (CI) growth using Graph Theory and causality analysis. Our findings highlight that, within the correlation networks of EV adoption, counties from the same state tend to cluster together, exhibiting greater homogeneity at finer time granularities. However, this pattern is not observed in CI growth networks. The homogeneity in clusters of counties for EV adoption suggests the importance of internal state-level characteristics that drive similar patterns of EV adoption within their counties.
Moreover, the correlations between EV adoption and CI growth are negligible in the two adoption scenarios: Early Adoption and Late Adoption. These correlations are very weakly positive, with slightly higher values in the latter scenario, where sufficient CI must be provided before EVs are widely adopted. This finding marginally favors the provision of sufficient CI as a strategy to promote EV adoption.
In our analysis of the causal effects between EV adoption and CI growth, we find a higher number of counties favoring Early Adoption compared to Late Adoption. This emphasizes the importance of policies and incentives aimed at boosting early EV adoption to spur subsequent investments in charging infrastructure. Additionally, our analysis of causal relationships shows that the causal effect of EV adoption on CI growth is slower than the reverse. Finally, 115 out of 137 counties demonstrate causal relationships under different combinations of time granularity, time lag, and adoption scenarios.
One potential direction for future work is to conduct an enrichment analysis of other factors, including demographic, economic, and political variables, for the counties that tend to cluster together. This would provide deeper insights into why these counties exhibit similar patterns in EV adoption. Another avenue for future research is to explore how different levels of CI relate to the two cases of Early and Late Adoption.